Please make sure your answer is correct. Don’t take the answers from previous question. If you are not sure leave the question for someone else.
media%2F07d%2F07db1f57-35a4-4fa8-9fb5-00
Answer #1

Equivalent unit = Unit transferred out+(ending work in process*Percent completion)

Material equivalent unit = 1130+(100*100%) = 1230 unit

Conversion equivalent unit = 1130+(100*50%) = 1180 unit